Suppose a consumer consuming two commodities X and Y has the following utility function X0.4 Y0.6. If price of good X and Y are 2 and 3 respectively and income constraints birr 50.
A/ Find the quantities of X and Y which Maximize utility
B/ Find the MRSxy
C/ What are the maximum utility drived from consuming of the two commodities?
1
Expert's answer
2020-01-24T09:46:37-0500

Utility functions U; X=0.4, Y=0.6 and prices P are 2 and 3 respectively. The income is 50

A. The quantity Q that of x and y that maximize utility is given by;

"Qx = Ux*Px" =0.4*2 =0.8

"Qy = Uy*Py" =0.6*3 =1.8

B. "MRSxy =Px\/Py"

MRSxy =2/3 =0.67

C. Maximum utility=0.8/0.67 +1.8/0.67=2.27